# DiagECG: An LLM-Driven Framework for Diagnostic Reasoning via Discretized ECG Tokenization

### v1

@@ Methods
All models based on large language models use LLaMA-3.2-3B as the default backbone, except ChatTime, which is built on LLaMA-2-7B. For methods involving LLM tuning, we apply 4-bit quantization with LoRA. LoRA rank is set to 64 with a scaling factor of 16. Models are trained for one epoch with a constant learning rate of 1e-4 and a batch size of 10. All experiments are conducted on one A100 GPU.
@@
Table 1 summarizes the performance of DiagECG and various baselines across ECG-QA and ECG-Report tasks on the MIMIC-IV-ECG and PTB-XL datasets. For ECG-QA, our model consistently achieves the highest exact match accuracy on both datasets, reaching 56.63% on MIMIC-IV-ECG and 54.98% on PTB-XL, with significant margins especially in the Query setting, which poses the greatest semantic challenge. This indicates that DiagECG can more effectively capture and interpret the clinical semantics of ECG signals compared to both TS-as-text and embedding-based LLM methods. In the ECG-Report task, DiagECG also obtains the best results across all evaluation metrics. On PTB-XL, it achieves a BLEU-4 score of 33.60, METEOR of 53.90, and ROUGE-L of 59.24, outperforming competitive models like TIME-LLM and LLaVA+W2V-CMSC-RLM. Our method also achieves strong performance on MIMIC-IV-ECG in a zero-shot setting, indicating robust generalization across datasets. These results validate the advantage of our discretization-based representation and token-level alignment between TS and language modalities.
@@
We conduct ablations on PTB-XL to isolate the effects of three core components in DiagECG. Results are shown in Figure 3. Removing the discretization module (w/o DISC) causes consistent drops across all tasks, suggesting that symbolic ECG tokens are more compatible with LLMs than raw continuous embeddings. Without fine-tuning the LLM (w/o FT), performance notably declines—especially on QA-Choose—indicating that LoRA-based adaptation helps the model capture task-specific semantics. Excluding tabular features (w/o TAB) slightly weakens generation and QA-Verify, reflecting their complementary role in providing patient context. Together, these results confirm that each component contributes meaningfully to downstream performance.
@@
We compare our lead-independent ResIncept encoder with four representative alternatives: MultiCNN (Oh et al. 2023), used in Fusion Transformer and MEIT; CausalCNN (Sun et al. 2024), the default encoder in TEST; Unet2D, a 2D convolution-based U-Net to extract channel-dependent features; and IndUnet1D, a structure that applies independent 1D U-Nets to each lead. As shown in Table 2, our encoder consistently outperforms all baselines across both datasets and all QA tasks. Notably, it yields the largest gains on QA-Query, which requires nuanced reasoning over subtle waveform features. Compared to MultiCNN and Unet2D, which either couple all leads early, our encoder better preserves fine-grained temporal features while maintaining lead-specific context. The improvement over CausalCNN suggests that fixed dilation patterns are less effective than our multi-scale design for ECG modeling.
Dataset Task Unet2D IndUnet1D MultiCNN CausalCNN Ours
MIMIC-
IV-ECG
QA-Verify 75.09 76.58 73.09 75.20 78.12
@@
Figure 4: Visualization of model alignment. (a) and (b) show attention maps for different clinical queries, with shaded areas denoting high-impact segments. (c) shows reconstructed waveforms from symbolic tokens, demonstrating information retention.
Assessing Latent Alignment Quality.
To assess whether the model captures clinically meaningful regions, we adopt a token-masking approach to estimate importance scores and visualize attention on the raw ECG waveform. Figure 4(a) shows that when asked about non-diagnostic T abnormalities, the model emphasizes the T-wave segments. In contrast, for the myocardial infarction query in Figure 4(b), attention shifts toward P and QRS complexes—consistent with diagnostic criteria. This context-dependent focus suggests that our alignment strategy successfully links ECG segments with query semantics. Additionally, Figure 4(c) confirms that the symbolic representation retains signal fidelity through accurate reconstruction.
@@
To examine the robustness of our learned representations, we evaluate zero-shot generalization by training on MIMIC-IV-ECG and directly testing on PTB-XL. As reported in Table 3, our method achieves the highest QA-Verify and QA-Query scores among all baselines, and the best average EM accuracy overall. Notably, although ST-MEM slightly outperforms our model on QA-Choose, its QA-Query performance remains weak. In contrast, our discretized token representation provides more balanced gains across all QA subtasks. These results indicate that symbolic representations learned from raw ECG signals promote better domain transfer and resist overfitting to dataset-specific artifacts.
@@ Conclusion
We present DiagECG, a unified framework that enables language-based clinical reasoning over ECG signals through symbolic tokenization and lightweight adaptation. By discretizing multi-lead ECG signals into a structured vocabulary, DiagECG bridges the modality gap between continuous physiological data and discrete language representations. It avoids reliance on explicit ECG-text alignment or contrastive supervision, instead leveraging autoregressive pretraining and instruction tuning to support open-ended tasks such as question answering and report generation. Extensive experiments across multiple benchmarks demonstrate that DiagECG achieves state-of-the-art performance and generalizes well to unseen ECG distributions. While effective, DiagECG assumes offline processing and does not support real-time ECG analysis. Future work may extend our approach to streaming settings and incorporate medical knowledge to improve interpretability.

### v2

@@ Methods
All models based on large language models use LLaMA-3.2-3B as the default backbone, except ChatTime, which is built on LLaMA-2-7B. For methods involving LLM tuning, we apply 4-bit quantization with LoRA. LoRA rank is set to 64 with a scaling factor of 16. Models are trained for one epoch with a constant learning rate of 1e-4 and a batch size of 10. For ECG discretization, the number of discrete levels is set to $K=6$ and the dimension of the quantized code to $D=4$, resulting in an ECG vocabulary of 1296 unique tokens. All experiments are conducted on one A100 GPU.
@@
Table 1 summarizes the performance of HeartLLM and various baselines across ECG-QA and ECG-Report tasks on the MIMIC-IV-ECG and PTB-XL datasets. For ECG-QA, our model consistently achieves the highest exact match accuracy on both datasets, reaching 56.63 on MIMIC-IV-ECG and 54.98 on PTB-XL, with significant margins especially in the Query setting, which poses the greatest semantic challenge. This indicates that HeartLLM can more effectively capture and interpret the clinical semantics of ECG signals compared to both TS-as-text and embedding-based LLM methods. In the ECG-Report task, HeartLLM also obtains the best results across all evaluation metrics. On PTB-XL, it achieves a BLEU-4 score of 33.60, METEOR of 53.90, and ROUGE-L of 59.24, outperforming competitive models like TIME-LLM and LLaVA+W2V-CMSC-RLM. Our method also achieves strong performance on MIMIC-IV-ECG in a zero-shot setting, indicating robust generalization across datasets. These results validate the advantage of our discretization-based representation and token-level alignment between TS and language modalities.
@@
We conduct ablations on PTB-XL to isolate the effects of three core components in HeartLLM. Results are shown in Figure 3. Removing the discretization module (w/o DISC) causes consistent drops across all tasks, suggesting that symbolic ECG tokens are more compatible with LLMs than continuous embeddings. Without fine-tuning the LLM (w/o FT), performance notably declines—especially on QA-Choose—indicating that LoRA-based adaptation helps the model capture task-specific semantics. Excluding tabular features in the prompt (w/o TAB) slightly weakens generation and QA-Verify, reflecting their complementary role in providing patient context. Together, these results confirm that each component contributes meaningfully to downstream performance.
@@
We compare our lead-independent ResIncept encoder with four representative alternatives: MultiCNN (Oh et al. 2023), used in Fusion Transformer and MEIT; CausalCNN (Sun et al. 2024), the default encoder in TEST; Unet2D, a 2D convolution-based U-Net to extract channel-dependent features; and IndUnet1D, a structure that applies independent 1D U-Nets to each lead. As shown in Table 2, our encoder consistently outperforms all baselines across both datasets and all QA tasks. Notably, it yields the largest gains on QA-Query, which requires nuanced reasoning over subtle waveform features. Compared to MultiCNN and Unet2D, which either couple all leads early, our encoder better preserves fine-grained temporal features while maintaining lead-specific context. The improvement over CausalCNN suggests that fixed dilation patterns are less effective than our multi-scale design for ECGs.
Dataset Task Unet2D Unet1D MultiCNN CausalCNN Ours
MIMIC- IV-ECG QA-Verify 75.09 76.58 73.09 75.20 78.12
@@
(a) PTB-XL Report
(b) PTB-XL QA-Verify
(c) PTB-XL QA-Choose
(d) PTB-XL QA-Query
Figure 3: Ablation study on the PTB-XL. (a) Report generation evaluated by BLEU, METEOR, and ROUGE. (b–d) EM accuracy on QA-Verify, QA-Choose, and QA-Query tasks. We compare the full model (Ours) against three variants: w/o TAB, w/o FT, and w/o DISC.
Figure 4: Visualization of latent alignment and signal reconstruction. (a-b) show attention maps for clinical queries, with shaded areas denoting high-impact segments. (c) shows reconstructed waveforms from symbolic tokens, demonstrating that discretization suppresses noise overfitting while preserving key information.
Evaluating Latent Alignment and Signal Reconstruction.
To assess whether the model captures clinically meaningful regions, we perform token-masking attribution: sequentially masking each ECG token and measuring the resulting drop in EM ACC to estimate its importance, which is then mapped back onto the original ECG waveform to visualize model attention. Figure 4(a) shows that when asked about non-diagnostic T abnormalities, the model emphasizes the T-wave segments. In contrast, for the myocardial infarction query in Figure 4(b), attention shifts toward P and QRS complexes—consistent with diagnostic criteria. This context-dependent focus suggests that our alignment strategy successfully links ECG segments with query semantics. Additionally, Figure 4(c) demonstrates that discretization acts as an implicit regularizer, suppressing overfitting to noisy ECG segments while preserving clinically informative waveform morphologies through accurate reconstruction.
@@
To examine the robustness of our learned representations, we evaluate zero-shot generalization by training on MIMIC-IV-ECG and directly testing on PTB-XL. As reported in Table 3, our method achieves the highest QA-Verify and QA-Query scores among all baselines, and the best average EM accuracy overall. Notably, although ST-MEM slightly outperforms our model on QA-Choose, its QA-Query performance remains weak. In contrast, our discretized token representation provides more balanced gains across all QA subtasks. These results indicate that the proposed symbolic ECG representations, implicitly aligned with the LLM’s textual embedding space, enhance domain transfer and resist overfitting to dataset-specific artifacts.
@@ Conclusion
We present HeartLLM, a unified framework that enables language-based clinical reasoning over ECG signals through symbolic tokenization and lightweight adaptation. By discretizing multi-lead ECG signals into a structured vocabulary, HeartLLM bridges the modality gap between continuous physiological data and discrete language representations. It avoids reliance on explicit ECG-text alignment or contrastive supervision, instead leveraging autoregressive pretraining and instruction tuning to support open-ended tasks such as question answering and report generation. Extensive experiments across multiple benchmarks demonstrate that HeartLLM achieves state-of-the-art performance and generalizes well to unseen ECG distributions. While effective, HeartLLM assumes offline processing and does not support real-time ECG analysis. Future work may extend our approach to streaming settings and incorporate medical knowledge to improve interpretability.